OpenStudy (brucebaner):

Which numbers are greater than 5 × 10−6? Choose all answers that are correct. A. 9 × 10−8 B. 7 × 10−6 C. 5 × 10−4 D. 4 × 10−6 E. 8 × 10−9

OpenStudy (anonymous):

Hi bruce baner :D In theory, you should be able to just complete the equation unless I'm missing something. In this case 5X10-6=44 A=9X10-8=82 B=7X10-6=64 C=5X10-4=46 Etc. Etc. Is this a bit clearer now?